# The MUSE eXtremely Deep Field: Detections of circumgalactic Siāii*Ā emission at ā ā thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

### v1

@@ Abstract
Context. The circumgalactic medium (CGM) serves as a baryon reservoir that connects galaxies to the intergalactic medium and fuels star formation. The spatial distribution of the metal-enriched cool CGM has not yet been directly revealed at cosmic noon (ā), as bright emission lines at these redshifts are not covered by optical integral field units.
Aims. To remedy this situation, we aim for the first-ever detections and exploration of extended Siāii*Ā emission (low-ionization state, LIS), referred to as āSiāii*Ā halosā, at redshifts ranging from to as a means to trace the metal-enriched cool CGM.
Methods. We use a sample of 39 galaxies with systemic redshifts of ā measured with the [Cāiii] doublet in the MUSE Hubble Ultra Deep Field catalog, which contains integration times spanning from to hours. We search for extended Siāii*1265, 1309, 1533 emission (fluorescent lines) around individual galaxies. We also stack a subsample of 14 UV-bright galaxies.
Results. We report five individual detections of Siāii* halos. We also confirm the presence of Siāii* halos in stacks for the subsample containing UV-bright sources. The other lines do not show secure detections of extended emission in either individual or stacking analyses. These detections may imply that the presence of metal-enriched CGM is a common characteristic for UV-bright galaxies. To investigate whether the origin of Siāii*Ā is continuum pumping as suggested in previous studies, we check the consistency of the equivalent width (EW) of Siāii*Ā emission and the EW of SiāiiĀ absorption for the individual halo object with the most reliable detection. We confirm the equivalence, suggesting that photon conservation works for this object and pointing toward continuum pumping as the source of Siāii*. We also investigate Siāii*Ā lines in a RAMSES-RT zoom-in simulation including continuum pumping and find ubiquitous presence of extended halos.
@@
Galaxies: high-redshift ā galaxies: formation ā galaxies: evolution ā galaxies: halos1 Introduction
@@
Historically, the CGM has been observed in absorption imprinted on quasi-stellar object (QSO) spectra (e.g., Wolfe etĀ al. 1986). This method can trace a wide range of column densities and ionization states of various elements. The CGM is found to be a multiphase medium in terms of its density, temperature, ionization, kinematics, metallicity, and structure (e.g., Werk etĀ al. 2016; Steidel etĀ al. 2016; Chen etĀ al. 2020; Schroetter etĀ al. 2021). However, this method is limited to the lines of sight of background sources. Large programs with recent or upcoming instruments targeting bright galaxies as background sources (tomographic mapping) can reach the IGM scale. For instance, The COSMOS Lyman-Alpha Mapping And Tomography Observations (CLAMATO) survey with Keck/LRIS (Lee etĀ al. 2014) and the Prime Focus Spectrograph-Subaru strategic program (PFS SSP) Galaxy Evolution survey (Greene etĀ al. 2022) have a few Mpc transverse resolution at . Higher spatial resolutions can be achieved using QSO pairs (e.g., Tytler etĀ al. 2009; UrbanoĀ Stawinski etĀ al. 2023), gravitationally lensed QSOs (e.g., Chen etĀ al. 2014; Rubin etĀ al. 2018), and gravitational-arcs as background sources (e.g., Lopez etĀ al. 2018). A galaxy-centered stacking approach is also useful to statistically map gas and metals with impact parameters ranging from kpc to a few Mpc (e.g., Rakic etĀ al. 2012; Turner etĀ al. 2014; Dutta etĀ al. 2024). Recently, Bordoloi etĀ al. (2023) achieves the impact parameters of ā kpc individually, using James Web Space Telescope (JWST)/NIRCam slitlesss grism spectroscopy. These studies enable to reach an outer-CGM scale, but an inner-CGM scale is difficult to investigate. Moreover, the identification of the host galaxies that are responsible for the absorption lines in QSO spectra is a challenging task. Hence, this method cannot provide the spatial distribution of gas and metals around individual host galaxies on an inner CGM scale.
@@
A complementary method to absorption-line studies is direct observations of the CGM in emission, which has been challenging because of its low surface brightness (SBs). Modern optical integral field units (IFUs) such as the Multi-Unit Spectroscopic Explorer (MUSE; Bacon etĀ al. 2010) and the Keck Cosmic Web Imager (KCWI; Morrissey 2018) make it possible to detect diffuse CGM emission around individual host galaxies (e.g., Wisotzki etĀ al. 2016). In particular, extended Ly emission, the so-called āLy haloā, which traces cool and warm hydrogen CGM (e.g., Steidel etĀ al. 2011; Momose etĀ al. 2014; Guo etĀ al. 2023c) has been intensively detected and studied around individual Ly emitters (LAEs) at ā (e.g., Leclercq etĀ al. 2017, 2020; Chen etĀ al. 2021; Claeyssens etĀ al. 2022; Erb etĀ al. 2023). Their diverse profiles and statistical trends as well as origins can now be discussed. Recently, a high fraction of Ly halos around UV-continuum selected (not based on Ly) galaxies at - is reported (Kusakabe etĀ al. 2022). The ubiquitous presence of resorvoirs of hydrogen in the CGM is directly confirmed at cosmic noon and earlier epochs ( with MUSE).
@@
However, metals in the CGM at cosmic noon (ā), which is the peak of star-forming activity of galaxies, have not been directly confirmed in emission, in particular for the cool phase. The CGM is known to be metal enriched even at ā from transverse absorption-line studies (e.g., Turner etĀ al. 2014; Lehner etĀ al. 2016; MĆ©ndez-HernĆ”ndez etĀ al. 2022; UrbanoĀ Stawinski etĀ al. 2023; Bordoloi etĀ al. 2023). This is also suggested by outflowing metals identified by down-the-barrel analysis (e.g., Shapley etĀ al. 2003; Steidel etĀ al. 2010; Du etĀ al. 2018; Sugahara etĀ al. 2017) and by broad-line components seen in spectra (e.g., Carniani etĀ al. 2023; Xu etĀ al. 2023). A popular emission tracer of cool metal-enriched CGM, MgāiiĀ 2796, 2803, is useful only at because at higher redshifts the observed wavelength shifts into the near-infrared and the cosmic dimming effect is more severe (e.g., Rubin etĀ al. 2011; Erb etĀ al. 2012; Martin etĀ al. 2013; Burchett etĀ al. 2021; Zabl etĀ al. 2021; Leclercq etĀ al. 2022; Dutta etĀ al. 2023; Guo etĀ al. 2023a, Pessa et al. 2024 in prep.). Other possible tracers are [Oāii] (e.g., Yuma etĀ al. 2013, 2017; Epinat etĀ al. 2018; Johnson etĀ al. 2018) and very faint Feāii*, , , and (e.g., Finley etĀ al. 2017; Shaban etĀ al. 2022). At ā, [Cāii]158m can be covered by sensitive ALMA bands 6 and 7 ([Cāii]Ā halos; e.g., Fujimoto etĀ al. 2019; Ginolfi etĀ al. 2020). At ā, no bright emission tracer of the cool metal-enriched CGM is available in the observed optical regime. While extremely faint, the most promising tracers are the emission lines Siāii*1265, 1309, 1533, which are not contaminated by other line features at nearby wavelengths given the spectral resolution of current facilities (cf. Cāii*1335 emission and Cāii1334 absorption, which cannot be resolved by the resolving power of MUSE, 3000).
@@
Siāii*Ā are fine-structure fluorescent emission lines from singly ionized silicon, which is a low-ionization state (LIS), with an ionization energy close to HāiĀ (ionization energy for Si and Si+ are 8.1Ā eV and 16.4Ā eV, respectively). The energy levels of the Si+ ion for the three transitions this paper focuses on are shown in Figure 1. The origin of Siāii*Ā is suggested to be ācontinuum pumpingā (or ācontinuum fluorescenceā) rather than collisional excitation or recombination (Shapley etĀ al. 2003). Collisional excitations are dominant only in dense environments with high electron densities, which is higher than typical values in Hāii regions (Shapley etĀ al. 2003). And the recombination rates of Si2+ into the excited Si+ is of the same order as the collisional excitation rates when Si+ and Si2+ have comparable abundance with K (Shull & van Steenberg 1982; Shapley etĀ al. 2003). Shapley etĀ al. (2003) model observed nebular emission lines using the photoionization code CLOUDY (Ferland etĀ al. 1998). They find that any model that reproduces the line ratios of the other lines ([Oāiii], [Oāii], H, Oāiii], and Cāiii]) predict more than one order of magnitude weaker Siāii*Ā emission by collisions and recombination. Therefore, those two processes are unlikely to be the dominant origin of Siāii*Ā emission in Hāii regions. As for the continuum pumping scenario, first, Si+ in neutral clouds absorbs the UV continuum and creates SiāiiĀ resonant absorption features. Second, a fraction of Si+ deexcites to the fine-structure level and emits Siāii*Ā photons (see Figure 1). Siāii*Ā lines are escape channels of resonant transitions and are expected to be as spatially extended as the resonant emission lines.
@@
Fluorescent lines are extremely faint, and Siāii*Ā emission is ā times weaker than Ly emission of LAEs (Steidel etĀ al. 2018). In fact, extended Siāii*Ā emission has not been directly detected. Wang etĀ al. (2020) found weak Siāii*Ā emission lines compared to the associated absorption line in HST/Cosmic Origins Spectrograph (COS) data for 5 galaxies at , which implies that the bulk of Siāii*Ā emission arises on larger scales than the COS aperture. Gazagnes etĀ al. (2023) compare COS spectra of Siāii*1265 for 45 local galaxies (COS Legacy Archive Spectroscopic SurveY, CLASSY) with mock spectra from zoom-in simulations (Mauerhofer etĀ al. 2021). They find stronger Siāii*Ā emission lines in the simulations than observed, and they argue that aperture losses of COS can explain the weakness of observed Siāii*Ā and that Siāii*1265 is spatially extended. Very recently, Vasan G.Ā C. etĀ al. (2024) study spatially-resolved outflow properties for a lensed star-forming galaxy at with KCWI and show that Siāii*Ā emission is more extended than the continuum for a certain direction extracted with a pseudo slit.
@@
In this project, we search for Siāii*Ā halos around galaxies at cosmic noon (at ā) as a tracer of metal-enriched cool CGM using integral field unit data to remedy the problem. We use data from the MUSE Hubble Ultra Deep Field (HUDF) survey (Bacon etĀ al. 2023) and test the presence of Siāii*Ā halos with surface brightness profiles of Siāii*Ā and UV continuum. We report the first detections of Siāii*Ā halos. We also stack the MUSE data for a UV-bright subsample to study the general extent of Siāii*Ā and compare it with simulations.
@@
The paper is organized as follows. In Section 2, we describe the data and the sample construction. Section 3 presents methods and results of the search for individual Siāii*Ā halos and those of a stacked subsample. In Section 4, we discuss photon conservation for Siāii, compare observed results with those of zoom-in simulations, and discuss how to increase a sample size with different selection criteria. Finally, conclusions are given in Section 5. Throughout this paper, we assume the Planck 2018 cosmological model (Planck Collaboration etĀ al. 2020) with a matter density of , a dark energy density of , and a Hubble constant of km s-1 Mpc-1 (). Magnitudes are given in the AB system (Oke & Gunn 1983). All distances are in physical units (kpc), unless otherwise stated.
@@ 2 Data and sample
To search for extended Siāii*Ā emission of galaxies, we utilize deep MUSE data from data release 2 (DR2) of the MUSE HUDF surveys (Bacon etĀ al. 2023), where rich multi-wavelength data are also available (see Section 2.1). We construct our sample with the DR2 catalog using the systemic redshifts for the individual search and stacking analyses (see Sections 2.2 and 2.3). We create masks to avoid potential contaminations from neighboring galaxies (see Section 2.4).
@@ 2.1 Data and catalogs
The MUSE HUDF DR2 data are obtained as a part of the MUSE guaranteed time observations (GTO) program described in Bacon etĀ al. (2023). We use two fields among the MUSE HUDF DR2 data set: a single arcmin2 pointing with a 31-hour depth (udf-10, see also Bacon etĀ al. 2017 and Inami etĀ al. 2017), and the adaptive-optics (AO) assisted MUSE eXtremely Deep Field (mxdf), whose maximum integration time is 141 hours with a 1-arcmin diameter field of view. They are located inside the Hubble eXtreme Deep Field (XDF; Illingworth etĀ al. 2013) with deep HST data (see Bacon etĀ al. 2023, for more details). The udf-10Ā covers the optical wavelength range from 4750Ā Ć Ā to 9350Ā Ć , while mxdfĀ covers from 4700Ā Ć Ā to 9350Ā Ć , with an AO gap from 5800Ā Ć Ā to 5966.25Ā Ć . The spectral resolving power of MUSE varies from to 3750 at 4700Ā Ć Ā to 9350Ā Ć , respectively. The full width at half maximums (FWHMs) of the Moffat point spread function (Moffat PSF, Moffat 1969) are () at the blue wavelength edge and () at the red wavelength edge for the mxdfĀ (udf-10). The point-source flux limit for an unresolved emission line in the mxdfĀ (the udf-10) is erg s-1 cm-2Ā ( erg s-1 cm-2), at around 7000Ā Ć Ā (not affected by OH sky emission), which corresponds to a surface brightness limit for an unresolved emission line of erg s-1 cm-2 arcsec-2Ā ( erg s-1 cm-2 arcsec-2, see Bacon etĀ al. 2023).
The MUSE HUDF DR2 catalog includes 2221 sources from to , which are selected by emission-line detections as well as through HST priors. The details of the catalog construction are described in Bacon etĀ al. (2023). Each MUSE object has a source file, which is in the format of MPDAF111https://mpdaf.readthedocs.io/en/latest/ multi-fits format and is composed of various data files such as minicubes, images, and spectra. It is available on the AMUSED website222https://amused.univ-lyon1.fr/. We use 5ā5ā minicubes from each of the source files. We create continuum-subtracted minicubes following the method used in Kusakabe etĀ al. (2022), taking the median for each pixel within a spectral window of 100 slices (masking km s-1 around the Ly wavelength if the spectrum covers). The continuum subtraction is useful not only to investigate emission lines, but also to remove neighboring sources around a target source. In order to identify HST counterparts, calculate UV magnitudes, and create masks, we use the HST catalog and data from Rafelski etĀ al. (2015).

### v2

# The MUSE eXtremely Deep Field: Detections of circumgalactic Si ii* emission at ††thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

@@ Abstract
Context. The circumgalactic medium (CGM) serves as a baryon reservoir that connects galaxies to the intergalactic medium and fuels star formation. The spatial distribution of the metal-enriched cool CGM has not yet been directly revealed at cosmic noon (–), as bright emission lines at these redshifts are not covered by optical integral field units.
Aims. To remedy this situation, we aim for the first-ever detections and exploration of extended Si ii* emission (low-ionization state, LIS), referred to as “Si ii* halos”, at redshifts ranging from to as a means to trace the metal-enriched cool CGM.
Methods. We use a sample of 39 galaxies with systemic redshifts of – measured with the [C iii] doublet in the MUSE Hubble Ultra Deep Field catalog, whose integration times span from to hours. We search for extended Si ii* 1265, 1309, 1533 emission (fluorescent lines) around individual galaxies. We also stack a subsample of 14 UV-bright galaxies.
Results. We report five individual detections of Si ii* halos. We also confirm the presence of Si ii* halos in stacks for the subsample containing UV-bright sources. The other lines do not show secure detections of extended emission in either individual or stacking analyses. These detections may imply that the presence of metal-enriched CGM is a common characteristic for UV-bright galaxies. To investigate whether the origin of Si ii* is continuum pumping as suggested in previous studies, we check the consistency of the equivalent width (EW) of Si ii* emission and the EW of Si ii absorption for the individual halo object with the most reliable detection. We confirm the equivalence, suggesting that photon conservation works for this object and pointing toward continuum pumping as the source of Si ii*. We also investigate Si ii* lines in a RAMSES-RT zoom-in simulation including continuum pumping and find ubiquitous presence of extended halos.
@@
Galaxies: high-redshift – galaxies: formation – galaxies: evolution – galaxies: halos1 Introduction
@@
Historically, the CGM has been observed in absorption imprinted on quasi-stellar object (QSO) spectra (e.g., Wolfe et al. 1986). This method can trace a wide range of column densities and ionization states of various elements. The CGM is found to be a multiphase medium in terms of its density, temperature, ionization, kinematics, metallicity, and structure (e.g., Werk et al. 2016; Steidel et al. 2016; Chen et al. 2020; Schroetter et al. 2021). However, this method is limited to the lines of sight of background sources. Large programs with recent or upcoming instruments targeting bright galaxies as background sources (tomographic mapping) can reach the IGM scale. For instance, The COSMOS Lyman-Alpha Mapping And Tomography Observations (CLAMATO) survey with Keck/LRIS (Lee et al. 2014) and the Prime Focus Spectrograph-Subaru strategic program (PFS SSP) Galaxy Evolution survey (Greene et al. 2022) have a few Mpc transverse resolution at . Higher spatial resolutions can be achieved using QSO pairs (e.g., Tytler et al. 2009; Urbano Stawinski et al. 2023), gravitationally lensed QSOs (e.g., Chen et al. 2014; Rubin et al. 2018), and gravitational-arcs as background sources (e.g., Lopez et al. 2018). A galaxy-centered stacking approach is also useful to statistically map gas and metals with impact parameters ranging from kpc to a few Mpc (e.g., Rakic et al. 2012; Turner et al. 2014; Dutta et al. 2024). Recently, bordoloi_eiger_2023 achieves the impact parameters of – kpc individually, using James Web Space Telescope (JWST)/NIRCam slitlesss grism spectroscopy. These studies enable to reach an outer-CGM scale, but an inner-CGM scale111Roughly within radii of , where is virial radius. The inner CGM is defined as kpc in Tumlinson et al. (2017) for L∗ galaxies at (with a median kpc in Werk et al. 2014). is difficult to investigate. Moreover, the identification of the host galaxies that are responsible for the absorption lines in QSO spectra is a challenging task. Hence, this method cannot provide the spatial distribution of gas and metals around individual host galaxies on an inner CGM scale.
@@
A complementary method to absorption-line studies is direct observations of the CGM in emission, which has been challenging because of its low surface brightness (SBs). Modern optical integral field units (IFUs) such as the Multi-Unit Spectroscopic Explorer (MUSE; Bacon et al. 2010) and the Keck Cosmic Web Imager (KCWI; Morrissey 2018) make it possible to detect diffuse CGM emission around individual host galaxies (e.g., Wisotzki et al. 2016). In particular, extended Ly emission, the so-called “Ly halo”, which traces cool and warm hydrogen CGM (e.g., Steidel et al. 2011; Momose et al. 2014; Guo et al. 2024) has been intensively detected and studied around individual Ly emitters (LAEs) at – (e.g., Leclercq et al. 2017, 2020; Chen et al. 2021; Claeyssens et al. 2022; Erb et al. 2023). Their diverse profiles and statistical trends as well as origins can now be discussed. Recently, a high fraction of Ly halos around UV-continuum selected (not based on Ly) galaxies at - is reported (Kusakabe et al. 2022). The ubiquitous presence of resorvoirs of hydrogen in the CGM is directly confirmed at cosmic noon and earlier epochs ( with MUSE).
@@
However, metals in the CGM at cosmic noon (–), which is the peak of star-forming activity of galaxies, have not been directly confirmed in emission, in particular for the cool phase. The CGM is known to be metal enriched even at – from transverse absorption-line studies (e.g., Turner et al. 2014; Lehner et al. 2016, 2022; Méndez-Hernández et al. 2022; Urbano Stawinski et al. 2023; Bordoloi et al. 2024; Banerjee et al. 2023; Beckett et al. 2024). This is also suggested by outflowing metals identified by down-the-barrel analysis (e.g., Shapley et al. 2003; Steidel et al. 2010; Du et al. 2018; Sugahara et al. 2017) and by broad-line components seen in spectra (e.g., Carniani et al. 2024; Xu et al. 2023). A popular emission tracer of cool metal-enriched CGM, Mg ii 2796, 2803, is useful only at because at higher redshifts the observed wavelength shifts into the near-infrared and the cosmic dimming effect is more severe (e.g., Rubin et al. 2011; Erb et al. 2012; Martin et al. 2013; Burchett et al. 2021; Zabl et al. 2021; Leclercq et al. 2022; Dutta et al. 2023; Guo et al. 2023a; Pessa et al. 2024). Other possible tracers are [O ii] (e.g., Yuma et al. 2013, 2017; Epinat et al. 2018; Johnson et al. 2018) and very faint Fe ii*, , , and (e.g., Finley et al. 2017; Shaban et al. 2022). At –, [C ii]158m can be covered by sensitive ALMA bands 6 and 7 ([C ii] halos; e.g., Fujimoto et al. 2019; Ginolfi et al. 2020). At –, no bright emission tracer of the cool metal-enriched CGM is available in the observed optical regime. While extremely faint, the most promising tracers are the emission lines Si ii* 1265, 1309, 1533, which are not contaminated by other line features at nearby wavelengths given the spectral resolution of current facilities (cf. C ii*1335 emission and C ii1334 absorption, which cannot be resolved by the resolving power of MUSE, – for –Å, or Å for C ii* line at ).
@@
Si ii* are fine-structure fluorescent emission lines from singly ionized silicon, which is a low-ionization state (LIS), with an ionization energy close to H i (ionization energy for Si and Si+ are 8.1 eV and 16.4 eV, respectively). The energy levels of the Si+ ion for the three transitions this paper focuses on are shown in Fig. 1. The origin of Si ii* is suggested to be “continuum pumping” (or “continuum fluorescence”) rather than collisional excitation or recombination (Shapley et al. 2003). Collisional excitations are dominant only in dense environments with high electron densities, which is higher than typical values in H ii regions (Shapley et al. 2003). And the recombination rates of Si2+ into the excited Si+ is of the same order as the collisional excitation rates when Si+ and Si2+ have comparable abundance with K (Shull & van Steenberg 1982; Shapley et al. 2003). Shapley et al. (2003) model observed nebular emission lines using the photoionization code CLOUDY (Ferland et al. 1998). They find that any model that reproduces the line ratios of the other lines ([O iii], [O ii], H, O iii], and C iii]) predicts more than one order of magnitude weaker Si ii* emission by collisions and recombination. Therefore, those two processes are unlikely to be the dominant origin of Si ii* emission in H ii regions. As for the continuum pumping scenario, first, Si+ in neutral clouds absorbs the UV continuum and creates Si ii resonant absorption features. Second, a fraction of Si+ deexcites to the fine-structure level and emits Si ii* photons (see Fig. 1). Si ii* lines are escape channels of resonant transitions and are expected to be as spatially extended as the resonant emission lines.
@@
Fluorescent lines are extremely faint, and Si ii* emission is – times weaker than Ly emission of LAEs (Steidel et al. 2018). In fact, extended Si ii* emission has not been directly detected. Wang et al. (2020) found weak Si ii* emission lines compared to the associated absorption line in HST/Cosmic Origins Spectrograph (COS) data for 5 galaxies at , which implies that the bulk of Si ii* emission arises on larger scales than the COS aperture. Gazagnes et al. (2023) compare COS spectra of Si ii* 1265 for 45 local galaxies (COS Legacy Archive Spectroscopic SurveY, CLASSY) with mock spectra from zoom-in simulations (Mauerhofer et al. 2021). They find stronger Si ii* emission lines in the simulations than observed, and they argue that aperture losses of COS can explain the weakness of observed Si ii* and that Si ii* 1265 is spatially extended. Very recently, Keerthi Vasan G. et al. (2024) study spatially-resolved outflow properties for a lensed star-forming galaxy at with KCWI and show that Si ii* emission is more extended than the continuum for a certain direction extracted with a pseudo slit.
@@
In this project, we search for Si ii* halos around galaxies at cosmic noon (at –) as a tracer of metal-enriched cool CGM using integral field unit data to remedy the problem. We use data from the MUSE Hubble Ultra Deep Field (HUDF) survey (Bacon et al. 2023) and test the presence of Si ii* halos with surface brightness profiles of Si ii* and UV continuum. We report the first detections of Si ii* halos. We also stack the MUSE data for a UV-bright subsample to study the general extent of Si ii* and compare it with simulations.
@@
The paper is organized as follows. In Section 2, we describe the data and the sample construction. Section 3 presents methods and results of the search for individual Si ii* halos and those of a stacked subsample. In Section 4, we discuss photon conservation for Si ii, compare observed results with those of zoom-in simulations, and discuss how to increase a sample size with different selection criteria. Finally, conclusions are given in Section 5. Throughout this paper, we assume the Planck 2018 cosmological model (Planck Collaboration et al. 2020) with a matter density of , a dark energy density of , and a Hubble constant of km s-1 Mpc-1 (). Magnitudes are given in the AB system (Oke & Gunn 1983). All distances are in physical units (kpc), unless otherwise stated.
@@ 2 Data and sample
To search for extended Si ii* emission of galaxies, we utilize deep MUSE data from data release 2 (DR2) of the MUSE HUDF surveys (Bacon et al. 2023), where rich multi-wavelength data are also available (see Section 2.1). We construct our sample with the DR2 catalog using the systemic redshifts for the individual search and stacking analyses (see Sections 2.2 and 2.3). We create masks to avoid potential contaminations from neighboring galaxies (see Section 2.4).
@@ 2.1 Data and catalogs
The MUSE HUDF DR2 data are obtained as a part of the MUSE guaranteed time observations (GTO) program described in Bacon et al. (2023). We use two fields among the MUSE HUDF DR2 data set: a single arcmin2 pointing with a 31-hour depth (udf-10, see also Bacon et al. 2017 and Inami et al. 2017), and the adaptive-optics (AO) assisted MUSE eXtremely Deep Field (mxdf), whose maximum integration time is 141 hours with a 1-arcmin diameter field of view. They are located inside the Hubble eXtreme Deep Field (XDF; Illingworth et al. 2013) with deep HST data (see Bacon et al. 2023, for more details). The mxdf is the deepest IFU survey ever performed, while the udf-10 is the second deepest available for this study, which also has multi-wavelength data and catalogs constructed in the same manner as mxdf. The udf-10 covers the optical wavelength range from 4750 Å to 9350 Å, while mxdf covers from 4700 Å to 9350 Å, with an AO gap from 5800 Å to 5966.25 Å. The spectral resolving power of MUSE varies from to 3750 at 4700 Å to 9350 Å, respectively. The full width at half maximums (FWHMs) of the Moffat point spread function (Moffat PSF, Moffat 1969) are () at the blue wavelength edge and () at the red wavelength edge for the mxdf (udf-10). The point-source flux limit for an unresolved emission line in the mxdf (the udf-10) is erg s-1 cm-2 ( erg s-1 cm-2), at around 7000 Å (not affected by OH sky emission), which corresponds to a surface brightness limit for an unresolved emission line of erg s-1 cm-2 arcsec-2 ( erg s-1 cm-2 arcsec-2, see Bacon et al. 2023).
The MUSE HUDF DR2 catalog includes 2221 sources from to , which are selected by emission-line detections as well as through HST priors. The details of the catalog construction are described in Bacon et al. (2023). Each MUSE object has a source file, which is in the format of MPDAF222https://mpdaf.readthedocs.io/en/latest/ multi-fits format and is composed of various data files such as minicubes, images, and spectra. It is available on the AMUSED website333https://amused.univ-lyon1.fr/. We use 5”5” minicubes from each of the source files. We create continuum-subtracted minicubes following the method used in Kusakabe et al. (2022), taking the median for each pixel within a spectral window of 100 slices (masking km s-1 around the Ly wavelength if the spectrum covers). The continuum subtraction is useful not only to investigate emission lines, but also to remove neighboring sources around a target source. In order to identify HST counterparts, calculate UV magnitudes, and create masks, we use the HST catalog and data from Rafelski et al. (2015).